OpenStudy (vikingjc13):

let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). −7 −3 10 33

MARC (marc_d):

\[=f((2)^2-6(2)-7)\]

MARC (marc_d):

\[=f(4-12-7)\]

MARC (marc_d):

\[=f(-15)\]

MARC (marc_d):

\[=-15+8\]

MARC (marc_d):

\[=?\]solve it @vikingjc13